OpenStudy (sleepyjess):

Use a calculator to find an approximate solution to the equation. Round any intermediate work and steps to three decimal places. Show all of your steps. 7ln (x + 2.8) = 11.2

OpenStudy (cwrw238):

ln(x +2.8) = 11.2/7n= 1.6 ln 4 = 1.386 ln 5 = 1.609 so x + 2.8 is close to but below 5

OpenStudy (cwrw238):

i suggest you start at 4.9 and plug in values 4.91, 4.92 etc to get a root to 2 decimals; - then sub in 3 decimal place values to finally get a root correct to 2 decimal places

OpenStudy (cwrw238):

try 4.92 : x-2.8 = ln 4.92 = 1.5933 try 2.94 x - 2.8 = ln 4.94 = 1.5973 try 4.96 = 1.6014 - looks like root is about 4.95 try 4.945 = 1.5983 - a little low 4.95 = 1.59938 - very close 4.955 = 1.6004 - this is higher tha 1.6 so the answer is 4.95 correct to 2 decimal places oh we need to subtract 2.8 root = 4.95-2.8 = 2.15 to 2 DP's

OpenStudy (cwrw238):

the correct answer to 3 decimal places is 2.153

OpenStudy (sleepyjess):

\(\Huge\frak\color{magenta}{Thank~You!}\)